# LOGen: Toward LiDAR Object Generation by Point Diffusion

### v2

@@ Appendix D Distribution of evaluated objects
In LiDAR datasets, the number of objects displays large variations among classes, not to mention the number of points in these objects depending on the classes, distance to the sensor and orientation. As an illustration, the classwise frequency of objects in nuScenes is pictured on Fig. 8 of this appendix. More statistics are available in [3].
traffic conebikem.cycleconstr. veh.trailercarbuspedestriantruckbarrier000.20.20.20.20.40.40.40.4Frequency
@@
Full statistics regarding the number of points and objects are available in Tab. 7 of this appendix.
LOGen 1-NNA $\downarrow$ COV $\uparrow$ FPD $\downarrow$ KPD  $\downarrow$ APC  $\uparrow$
[…]
Table 6: Single-model-per-class vs. multi-class model. Comparison of a same size single-model-per-class and multi-class models on generation metrics. The multi-class model lags behind the single-model-per-class approach in realism metrics.
[…]
@@
Appendix E One multi-class vs Single model per class
Training single model per class allows us to focus on specific classes, maintain consistent batch sizes with minimal zero-padding, keep the network compact (7.5M parameters), and reduce overall computational overhead (measured in iterations per instance). However, a multi-class model of the same size and trained with the same number of iterations does not perform as well as a single model per class, as shown in Table 6. Notably, training single model per class introduces no drawbacks in terms of runtime or model efficiency. We note that DiT3D [27] similarly provides mostly class-wise results.
@@ Appendix F Detailed results
F.1 Classwise numbers
@@
LOGen PE 4 ch 0.08 0.12 0.15 0.13 0.15 0.11 0.11 0.09 0.2 0.16 0.130
EMD MinkUNet 0.16 0.22 0.2 0.13 0.2 0.19 0.27 0.35 0.24 0.16 0.212
DiT-3DL VE 3 ch 0.07 0.1 0.2 0.11 0.16 0.08 0.08 0.06 0.2 0.15 0.121
DiT-3DL PE 4 ch 0.06 0.09 0.21 0.11 0.15 0.08 0.1 0.07 0.21 0.17 0.125
PixArt-L PE 4 ch 0.1 0.13 0.39 0.3 0.22 0.14 0.1 0.1 0.29 0.27 0.204
@@
Table 8: Point-based metrics. [Complement to Table 1 in the paper] We present class-wise results of CD and EMD.
@@
Table 9: Point-based metrics. [Complement to Table 1 in the paper] We present class-wise results of 1-NNA.
@@
LOGen PE 4 ch 44.45 42.3 40.4 44.2 42.1 46.7 46.0 41.3 32.8 41.7 42.3
Table 10: Point-based metrics. [Complement to Table 1 in the paper] We present class-wise results of COV.
@@
Table 11: Point-based metrics. [Complement to Table 1] We present class-wise results of COV and 1-NNA on intensity realism.
@@
Table 12: Feature-based metrics. [Complement to Table 2 in the paper] We present class-wise results of FPD.
@@
Table 13: Feature-based metrics. [Complement to Table 2 in the paper] We present class-wise results of KPD.
@@
Table 14: Feature-based metrics. [Complement to Table 2 in the paper] We present class-wise results of APC.
3 ch. $d$ 0.37 1.78 1.44 0.60 1.49 0.79 0.14 1.52 4.16 1.02 1.34
$d\mathbin{/}2$ 0.44 4.95 2.75 2.72 3.22 1.72 1.13 4.16 5.80 3.94 3.08
$d\times 1.5$ 1.19 1.13 1.46 0.85 2.19 1.45 3.64 0.77 6.19 1.23 2.01
$d\times 2$ 2.80 4.12 2.98 1.80 3.63 3.03 13.53 5.15 8.71 3.63 4.79
4 ch. $d$ 0.64 3.78 1.89 0.56 3.04 3.03 0.52 0.99 4.93 2.31 2.18
$d\mathbin{/}2$ 2.17 9.92 11.88 5.40 9.80 7.01 4.27 4.73 16.68 9.44 8.13
$d\times 1.5$ 1.57 4.48 6.20 4.64 7.35 6.70 8.08 5.65 5.48 6.04 5.62
$d\times 2$ 3.38 11.67 12.23 12.82 13.03 10.64 22.78 17.89 10.28 12.29 12.7
Table 15: Evaluation of conditioning on distance to sensor. [Complement to Table 3 in the paper] We present class-wise results of FPD.
3 ch. $d$ 0.00 0.01 0.23 0.04 0.00 -0.00 -0.01 0.22 0.31 0.22 0.10
$d\mathbin{/}2$ 0.08 0.54 0.13 0.32 0.20 0.14 0.23 0.81 0.48 0.39 0.33
$d\times 1.5$ 0.13 0.05 0.06 0.02 0.10 0.05 0.67 0.04 0.58 0.06 0.18
$d\times 2$ 0.43 0.41 0.16 0.14 0.34 0.00 2.31 0.98 1.01 0.10 0.61
4 ch. $d$ 0.03 0.14 0.11 0.04 0.09 0.02 0.06 0.04 0.48 0.17 0.12
$d\mathbin{/}2$ 0.08 1.00 1.09 0.74 0.88 0.36 0.73 0.57 2.17 0.91 0.85
$d\times 1.5$ 0.10 0.57 0.63 0.59 0.65 0.59 1.33 0.83 0.37 0.74 0.64
$d\times 2$ 0.42 1.44 1.31 1.92 1.48 1.25 3.66 3.10 0.83 1.69 1.71
Table 16: Evaluation of conditioning on distance to sensor. [Complement to Table 3 in the paper] We present class-wise results of KPD.
3 ch. $d$ 65.7 6.3 7.8 49.4 30.8 19.8 86.3 94.1 32.0 16.1 40.9
$d\mathbin{/}2$ 67.1 4.9 11.7 46.3 40.1 22.7 90.1 84.5 27.7 17.0 38.3
$d\times 1.5$ 63.6 5.4 5.4 51.1 22.7 18.9 75.5 87.9 33.2 14.5 37.8
$d\times 2$ 60.9 3.2 3.6 53.2 14.3 19.1 66.9 73.7 32.1 12.5 34.0
4 ch. $d$ 72.9 8.6 9.6 58.5 49.0 24.5 97.7 94.3 49.9 16.4 48.1
$d\mathbin{/}2$ 72.8 8.3 17.0 63.0 49.5 23.0 97.7 91.3 56.0 19.4 49.8
$d\times 1.5$ 72.2 8.3 5.1 44.8 39.9 21.8 90.1 82.3 49.0 14.5 42.8
$d\times 2$ 71.1 6.0 4.1 35.1 32.0 22.7 81.2 68.7 49.1 13.8 38.4
Table 17: Evaluation of conditioning on distance to sensor. [Complement to Table 3 in the paper] We present class-wise results of APC.
Test Objects Box Orientation bar. bic. bus car c.veh mot. ped. t.con. tra. tru. mean
3 channel real orig. 80.3 48.8 90.7 94.2 40.6 86.0 085.5 73.4 60.7 85.2 74.5
LOGen orig. 78.2 50.9 60.1 79.1 28.1 82.2 84.2 73.5 37.5 63.2 63.7
4 channel real orig. 81.6 50.4 90.9 94.6 44.5 88.1 85.6 77.4 64.6 85.6 76.3
LOGen orig. 78.0 48.9 71.7 91.3 37.1 82.0 84.8 79.1 45.1 72.9 69.1
Table 18: Evaluation of conditioning on sensor viewing angle. [Complement to Table 5 in the paper] We present per class mIoU performances.
We provide complementary class-wise results for all our experiments in the main paper. For additional results on point-based metrics, refer to Table 8, Table 9, Table 10, and Table 11. For feature-based metrics, see Table 12, Table 13, and Table 14. Finally, for complementary results on the evaluation of conditioning on distance to sensor refer to Table 15, Table 16, Table 17, and evaluation of conditioning on sensor viewing angle is presented in Table 18.
F.2 More qualitative examples
In Figure 9, we present novel objects produced by LOGen for the remaining classes. To get a better sense of our 3D generations, please refer to the accompanying video results.
barrierp
bikep
busp
carp
[…]
constr. veh.p
motorcyclep
pedestrianp
traffic conep
trailerp
truckp

### v3

@@ Appendix D Distribution of evaluated objects
In LiDAR datasets, the number of objects displays large variations among classes, not to mention the number of points in these objects depending on the classes, distance to the sensor and orientation. As an illustration, the classwise frequency of objects in nuScenes is pictured on Fig. 7. More statistics are in [5] and Table 8.
Appendix E Evaluation of conditioning
Evaluation of conditioning on sensor viewing angle
To assess LOGen’s ability to generate realistic objects regardless of viewing angle, we modify the synthetic replacements by rotating their angular condition. This reduces the need to find suitable locations for new objects. We consider five evenly spaced angles around a full turn, i.e., $\nicefrac{{i}}{{5}}\times 360$° where $i\in\{0..4\}$, as shown in Figure 1.
Tab. 6 shows a moderate variation of the performance when rotating objects. It is partly due to collisions between objects that occur after rotation (the insertion is performed without such a test). Further, the distribution of object viewing angles in the dataset is not uniform. Less viewed angles can result in noisier generations.
traffic conebikem.cycleconstr. veh.trailercarbuspedestriantruckbarrier00.20.20.40.4Frequency
@@ Appendix E Evaluation of conditioning
Test Box mIoU (%) $\uparrow$
objects orientation 3 ch. 4 ch.
real orig. 74.5 76.3
LOGen orig. 63.7 69.1
LOGen rot. $\nicefrac{{1}}{{5}}$ 53.7 64.6
LOGen rot. $\nicefrac{{2}}{{5}}$ 54.8 66.0
LOGen rot. $\nicefrac{{3}}{{5}}$ 53.8 63.8
LOGen rot. $\nicefrac{{4}}{{5}}$ 53.3 64.1
Table 6: Evaluation of conditioning parameters: sensor viewing angle. As in Tab. 4 of Sec. 4.4, 50% of real objects are replaced by generated objects, and we evaluate semantics segmentation (mIoU%). Here we generate new objects with a given rotation relatively to the original viewing angle.
Distance FPD $\downarrow$ KPD $\downarrow$ APC (%) $\uparrow$
to sensor 3 ch. 4 ch. 3 ch. 4 ch. 3 ch. 4 ch.
orig. $d$ 1.34 2.18 0.10 0.12 40.9 48.1
$d\times 2$ 4.79 12.70 0.61 1.71 34.0 38.4
$d\mathbin{/}2$ 3.08 8.13 0.33 0.85 38.3 49.8
Table 7: Evaluation of conditioning parameters: distance to the sensor. We vary the distance when generating objects.
obj. size |s|<|s|\,{<} 20 pts objects size $|s|\,{\geq}$ 20 points
@@
Test objects Box orient. bar. bic. bus car c.veh mot. ped. t.con. tra. tru. mean
3 ch. real orig. 80.3 48.8 90.7 94.2 40.6 86.0 085.5 73.4 60.7 85.2 74.5
LOGen orig. 78.2 50.9 60.1 79.1 28.1 82.2 84.2 73.5 37.5 63.2 63.7
4 ch. real orig. 81.6 50.4 90.9 94.6 44.5 88.1 85.6 77.4 64.6 85.6 76.3
LOGen orig. 78.0 48.9 71.7 91.3 37.1 82.0 84.8 79.1 45.1 72.9 69.1
Table 9: Evaluation of conditioning: sensor viewing angle. We present per-class IoU% performances.
Evaluation of conditioning on distance to sensor.
@@ Appendix F Detailed results
Class-wise numbers on nuScenes.
Statistics on the number of points and objects in nuScenes are in Table 8.
For our experiments on nuScenes, we provide here results per class: barrier (bar.), bicycle (bic.), bus, car, construction vehicle (c.veh), motorbike (mot.), pedestrian (ped.), traffic cone (t.con.), trailer (tra.), truck (tru.).
Complementary class-wise results regarding the evaluation of conditioning on sensor viewing angle are shown in Table 9. Results regarding the conditioning on distance to sensor are in Table 11, Table 12, Table 13.
For class-wise results on point-based metrics, refer to Table 14, Table 15, Table 16, and Table 17. For feature-based metrics, see Table 18, Table 19, and Table 20.
Class CD $\downarrow$ EMD $\downarrow$ JSD $\downarrow$ FPD $\downarrow$ KPD $\downarrow$
bicycle 0.28 0.30 0.24 3.00 0.03
car 0.29 0.39 0.17 1.00 0.21
person 0.14 0.11 0.18 1.29 0.11
mean 0.24 0.27 0.19 1.76 0.12
Table 10: Class-wise results of LOGen on KITTI-360.
Class-wise numbers on KITTI-360.
In Table 10, we also provide class-wise results of KITTI-360, along with metrics CD, EMD, JSD, FPD and KPD.
3 ch. $d$ 0.37 1.78 1.44 0.60 1.49 0.79 0.14 1.52 4.16 1.02 1.34
$d\mathbin{/}2$ 0.44 4.95 2.75 2.72 3.22 1.72 1.13 4.16 5.80 3.94 3.08
$d\times 1.5$ 1.19 1.13 1.46 0.85 2.19 1.45 3.64 0.77 6.19 1.23 2.01
$d\times 2$ 2.80 4.12 2.98 1.80 3.63 3.03 13.53 5.15 8.71 3.63 4.79
4 ch. $d$ 0.64 3.78 1.89 0.56 3.04 3.03 0.52 0.99 4.93 2.31 2.18
$d\mathbin{/}2$ 2.17 9.92 11.88 5.40 9.80 7.01 4.27 4.73 16.68 9.44 8.13
$d\times 1.5$ 1.57 4.48 6.20 4.64 7.35 6.70 8.08 5.65 5.48 6.04 5.62
$d\times 2$ 3.38 11.67 12.23 12.82 13.03 10.64 22.78 17.89 10.28 12.29 12.70
Table 11: Evaluation of conditioning on distance to sensor: class-wise FPD.
3 ch. $d$ 0.00 0.01 0.23 0.04 0.00 0.00 -0.01 0.22 0.31 0.22 0.10
$d\mathbin{/}2$ 0.08 0.54 0.13 0.32 0.20 0.14 0.23 0.81 0.48 0.39 0.33
$d\times 1.5$ 0.13 0.05 0.06 0.02 0.10 0.05 0.67 0.04 0.58 0.06 0.18
$d\times 2$ 0.43 0.41 0.16 0.14 0.34 0.00 2.31 0.98 1.01 0.10 0.61
4 ch. $d$ 0.03 0.14 0.11 0.04 0.09 0.02 0.06 0.04 0.48 0.17 0.12
$d\mathbin{/}2$ 0.08 1.00 1.09 0.74 0.88 0.36 0.73 0.57 2.17 0.91 0.85
$d\times 1.5$ 0.10 0.57 0.63 0.59 0.65 0.59 1.33 0.83 0.37 0.74 0.64
$d\times 2$ 0.42 1.44 1.31 1.92 1.48 1.25 3.66 3.10 0.83 1.69 1.71
Table 12: Evaluation of conditioning on distance to sensor: class-wise KPD.
Note: KPD computation (like KID) uses an unbiased U-statistic estimator of MMD2 from finite samples. While its expected value equals the true (nonnegative) MMD2, individual estimates can dip below zero due to sampling noise. This explains why some values in the table are “slightly” negative.
3 ch. $d$ 65.7 6.3 7.8 49.4 30.8 19.8 86.3 94.1 32.0 16.1 40.9
$d\mathbin{/}2$ 67.1 4.9 11.7 46.3 40.1 22.7 90.1 84.5 27.7 17.0 38.3
$d\times 1.5$ 63.6 5.4 5.4 51.1 22.7 18.9 75.5 87.9 33.2 14.5 37.8
$d\times 2$ 60.9 3.2 3.6 53.2 14.3 19.1 66.9 73.7 32.1 12.5 34.0
4 ch. $d$ 72.9 8.6 9.6 58.5 49.0 24.5 97.7 94.3 49.9 16.4 48.1
$d\mathbin{/}2$ 72.8 8.3 17.0 63.0 49.5 23.0 97.7 91.3 56.0 19.4 49.8
$d\times 1.5$ 72.2 8.3 5.1 44.8 39.9 21.8 90.1 82.3 49.0 14.5 42.8
$d\times 2$ 71.1 6.0 4.1 35.1 32.0 22.7 81.2 68.7 49.1 13.8 38.4
Table 13: Evaluation of conditioning on distance to sensor: class-wise APC (%).
@@
LOGen PE 4 ch 0.08 0.12 0.15 0.13 0.15 0.11 0.11 0.09 0.20 0.16 0.130
EMD MinkUNet 0.16 0.22 0.20 0.13 0.20 0.19 0.27 0.35 0.24 0.16 0.212
DiT-3DL VE 3 ch 0.07 0.10 0.20 0.11 0.16 0.08 0.08 0.06 0.20 0.15 0.121
DiT-3DL PE 4 ch 0.06 0.09 0.21 0.11 0.15 0.08 0.10 0.07 0.21 0.17 0.125
PixArt-L PE 4 ch 0.10 0.13 0.39 0.3 0.22 0.14 0.10 0.10 0.29 0.27 0.204
@@
Table 14: Class-wise distance metrics CD and EMD.
@@
Table 15: Class-wise 1-NNA (%) metric using either CD or EMD as distance.
@@
LOGen PE 4 ch 44.5 42.3 40.4 44.2 42.1 46.7 46.0 41.3 32.8 41.7 42.3
Table 16: Class-wise COV (%) metric using either CD or EMD as distance.
@@
Table 17: Point-based metrics. We present class-wise results of COV (%) and 1-NNA (%) on intensity realism.
@@
Table 18: Feature-based metrics. We present class-wise results of FPD.
@@
Table 19: Feature-based metrics. We present class-wise results of KPD. (See also caption of Tab. 12.)
@@
Table 20: Feature-based metrics. We present class-wise results of APC.